a reaction of 12.640g of copper oxide with 0.316g hydrogen produces 2.844g of water and x g of copper. x = ____ g
1
Expert's answer
2020-09-30T04:17:20-0400

Question #135962

a reaction of 12.640g of copper oxide with 0.316g hydrogen produces 2.844g of water and x g of copper. x = ____ g

Answer

CuO (s) +H2(g) –> Cu(s)+H2O(g)

Mol of CuO= 12.64/80=.0158

Mol of H2 = 0.316/2=0.158

Mol of H2O = 2.844/18= 0.158

Mol ratio of CuO: H2: Cu: H2O is 1:1:1:1

Mol of Cu = 1×0.158=0.158

Mass of Cu formed, X = 0.158×64

                                  = 10.112g